Let f(x) = x2 and g(x) = x+1

____

x+3

1. Evaluate(f °g)(2)

2. Find formular for (f ° g) (x)


1
Expert's answer
2021-10-12T09:15:29-0400

1.


"(f\circ g)(x)=(\dfrac{x+1}{x+2})^2"

"(f\circ g)(2)=(\dfrac{2+1}{2+2})^2=\dfrac{9}{16}"

2.


"(f\circ g)(x)=(\dfrac{x+1}{x+2})^2=\dfrac{(x+1)^2}{(x+2)^2}, x\not=-2"